# SoK: LLM-based Log Parsing

### v1

@@ 5.3. Evaluation Metrics
Following previous work (Jiang et al., 2024b) we select Group Accuracy (GA), Parsing Accuracy (PA), and F1-score of Template Accuracy (FTA) which are described in Sec. 3.2. There, we also described the Edit Distance (ED) from which we compute the Normalized Edit Distance (NED) (Marzal and Vidal, 1993) following (Xiao et al., 2024).
GA and PA are sensitive to the total number of log messages, which is problematic since most log datasets contain a large number of logs but a much smaller number of unique templates (Jiang et al., 2024b). Therefore, Jiang et al. (Jiang et al., 2024b) proposed to use F-metrics since they are insensitive to class imbalance which is why we use FTA. Furthermore, we use NED since it constitutes a similarity metric between groundtruth and predicted template. The normalization makes NED also a more intuitive indicator for the effectiveness than ED.
@@ 5.4.1. Settings
The evaluation was performed using three different LLMs of different sizes (small, medium, large):
@@
    GPT-3.5131313https://platform.openai.com/docs/models/gpt-3.5-turbo; accessed 7-March-2025 (gpt-3.5-turbo-0125) is a widely known closed-source model, reaching outstanding performance in the earlier stages of the LLM hype. It was chosen because it is the most used one from all reviewed papers and it is considered to be medium sized (the size is actually undisclosed but a paper by researchers of Microsoft (Singh et al., 2023) state it is 20 billion parameters). The model is proprietary and was accessed via the OpenAI API.
@@
    DeepSeek R1141414https://github.com/deepseek-ai/DeepSeek-R1; accessed 25-February-2025 utilizes a novel reinforcement learning approach for model training. It has recently received wide public attention due to its outstanding performance, while being relatively small (but still large) compared to similarly performing models like OpenAI’s o1 model, in combination with its open-source release. This model is used to assess the potential gains in accuracy that can be achieved by utilizing one of the most effective models (671 billion parameters). DeepSeek R1 was called via API from TogetherAI [^15] https://www.together.ai/; accessed 14-March-2025.
The runtime evaluation with the LogHub-2.0 dataset (Jiang et al., 2024b) was run with GPT-3.5 on an Ubuntu 24.04 LTS server with a 32-core AMD EPYC-Milan CPU and 64 GB RAM.
@@ 5.4.2. Configuration and Code Changes
For all parsers we set the initial temperature of the LLMs to 00 for a more deterministic output. Note that some parsers increase the temperature in case of a template mismatch or other reasons for some of their LLM calls.
@@
As mentioned in Sec. 4.13.1, there are some issues with some of the parsers’ code which had to be resolved for the benchmark. To ensure the comparability of the output of the parsers, other changes were also necessary to the code of each parser. These changes are kept as small as possible to not disturb the original design or interfere with performance, or at least as little as possible. To ensure model or API compatibility, we updated deprecated versions of the OpenAI (python) package to the newer working ones and we added the functionality to call the OpenAI, Ollama, and TogetherAI API if not given. To have consistent output format, we ensure that the template placeholder symbols are always <*> and that the output is always a CSV file with at least the columns “Content”, and “EventTemplate”. “Content” refers to the log filtered by the log format (i.e. the log message). Since some parsers do not only parse the log content but the entire log, we modify the input to these parsers to only parse the content as well. As found by He et al. (He et al., 2016b), this typically improves the parsing accuracy.
@@
Specific changes were made to DivLog (Xu et al., 2024) which is one of the earliest works on LLM-based log parsing. It does not feature a caching mechanism and therefore, calls the LLM for each log line. Since this is a significant burden in terms of runtime (and cost) and therefore also unrealistic for real-life adaptation, we added a simple cache consisting only of the set of already parsed templates. The cache returns the corresponding template in the event of a match instead of calling the LLM. This slightly affects the accuracy metrics due to improved output consistency. Additionally, we implemented a minor post-processing step where multiple consecutive wildcards were replaced by a single wildcard <*> for efficient matching, to resolve an issue with the generation of infinite consecutive wildcards for some of the logs. This measure slightly increases DivLog’s accuracy.
@@
LogPrompt (Liu et al., 2024) accepts a maximum prompt size parameter which is by default $3000$, corresponding to roughly $25$ logs parsed at once (as they state). Early experiments revealed that this performs extremely poor, hence we set the maximum prompt size to $1000$. This corresponds to roughly $5$ to $10$ logs per prompt, which is in the recommended range of simultaneously parsed logs by Xiao et al. (Xiao et al., 2024). If a template cannot be matched with a log it is again sent to the LLM. We limit the maximum number of re-prompts to $3$ to avoid infinite LLM calls. LogPrompt also required some code corrections, other than typos. For some datasets, LogPrompt runs into an infinite loop due to a faulty template extraction process from the LLM responses. This was fixed by computing the enumeration of the logs instead of extracting it from the prompt again. From the paper it is not clear whether they used ICL, CoT or the self-prompt paradigm for their parsing performance evaluation, yet the parsing functionality is only given for the self-prompt case.
@@ 5.4.3. Sampling for Supervised Parsers
From the selected parsers LILAC (Jiang et al., 2024a) and DivLog (Xu et al., 2024) are supervised parsers and require labeled logs to achieve competitive accuracy. In this study, we emphasize minimal human effort and therefore keep the number of available templates $n$ small, namely $n=2$ and $n=4$. Note that the Apache dataset from LogHub (Zhu et al., 2023) contains the smallest number of templates, which is only $6$. For LILAC and DivLog it has been shown that the performance increases with an increasing number of labeled logs available (Jiang et al., 2024a; Xu et al., 2024). This is obvious as the similarity search retrieves the exact template the target log belongs to with a higher likeliness. Multiple works from our selection (Jiang et al., 2024a; Xu et al., 2024) as well as (Le and Zhang, 2023b) state that having a diverse candidate set is crucial to reduce the risk of overfitting to a specific log template, which is why they create specific sampling algorithms to sample from labeled logs. However, having a large proportion of templates at hand is an unrealistic setting in real-life application and requires users to label logs manually (Huang et al., 2024).
@@ 6.1.1. Baseline Performance
The performance of the baseline parsers is visualized in Fig. 3. For some reason, the implementation of ULP (Sedki et al., 2022) outputs the templates in a special format which is why the matching process of the evaluation function does not recognize correctly parsed logs, thus PA and FTA are 00 (GA is not affected).
@@ 6.1.2. Performance of LLM-based Parsers
The evaluation results of the parsers with the three LLMs on LogHub (Zhu et al., 2023) are given in Fig. 4(a), Fig. 4(b), and Fig. 4(c). Note that we excluded LogPrompt from the evaluation with DeepSeek R1 because the LLM was barely able to extract any templates, resulting in scores of roughly 00 in experimental runs. The LLM did not provide the output in the structured format that was requested. Therefore, the template extraction process, which is based on markers positioned directly before the extracted template, was not able to properly extract the template. This resulted in long inner monologues of DeepSeek R1 and maximum reprompts due to log-template mismatches which would have resulted in an unreasonable financial expense given the cost of the LLM and the poor performance.
@@
In the boxplots of Fig. 4(a), Fig. 4(b), and Fig. 4(c) we can see that the best performance on each metric and for each LLM is achieved by LogBatcher (Xiao et al., 2024), followed by LILAC (Jiang et al., 2024a) for $2$ and $4$ shots. LogBatcher and LILAC clearly outperform the conventional parsers, while the conventional parsers outperform the rest of the LLM-based parsers. In general, the performance of each LLM parser is rather stable across the different models. CodeLlama visibly performs worst, but it is also by far the smallest of the models with 7 billion parameters. Given the size difference of GPT-3.5 and DeepSeek R1, it is surprising that it does not outperform GPT-3.5, given its size. Furthermore, we report the exact numbers for the evaluation with GPT-3.5 in Table 5 to show how the parsers performed on each individual dataset.
Table 5. Performance of the selected parsers with GPT-3.5 on the corrected LogHub datasets including Audit
@@ 6.1.3. LLM General Performance
Figure 5. Averaged performance of all parsers for CodeLlama, GPT-3.5 and DeepSeek R1.
@@ 6.1.5. Performance on the Audit Dataset
An analysis of the performance of all parsers on the Audit dataset reveals that only the supervised parser, LILAC (Jiang et al., 2024a), and to some extent also DivLog (Xu et al., 2024), are able to parse the logs effectively. It should be noted that the configuration of the baseline parsers for Audit was not specifically tuned. It appears that the intended templates’ format of the logs is not straightforward for GPT-3.5 without supervised demonstrations. This finding underscores the efficacy of supervised parsers, demonstrating that a mere two templates are sufficient to attain a reasonable performance level, potentially even a single one demonstrating the preferred template style. Nevertheless, it should be noted that this is merely a solitary example. Future research could establish a more extensive benchmark by incorporating a range of novel and diverse datasets.
Figure 7. Performance of the selected parsers with GPT-3.5 only on the Audit data.
@@ 6.2. Efficiency
Figure 8 visualizes the computation time and LLM invocation time of the parsers, Fig. 9 shows the number of LLM calls made. The invocation time for conventional parsers is, of course, zero. Outstanding runtime efficiency is achieved by the conventional parser ULP (Sedki et al., 2022) on both datasets. While LLM-TD is the second fastest and also calls the LLM the least times it is rather the opposite on the BGL dataset. As mentioned in Sec. 4.13.1 and 5.4.2 there were some implementation problems with SelfLog (Pei et al., 2024). They provided a faulty script for the efficiency evaluation which is why we ran the evaluation with the script they designed for the evaluation of the $2000$ log lines version of LogHub (Zhu et al., 2023). However, for both HDFS and BGL the processes got killed and are therefore not featured in the plots.
@@ 7.1. Answers to Research Questions
LLM-based log parsing approaches offer key advantages such as adaptability to diverse log formats, the ability to generalize across different system logs, and robustness to unseen log templates as demonstrated by the evaluation on the Audit dataset (Landauer et al., 2023b, 2021) and other work (Fu et al., 2022). Approaches like the unsupervised parser LogBatcher (Xiao et al., 2024) and the supervised parser LILAC (Jiang et al., 2024a) can handle complex and unstructured logs more effectively than traditional methods. However, they also come with drawbacks, including higher computational costs, potential hallucinations and overthinking for reasoning models (Cuadron et al., 2025), and difficulty in interpreting results due to their black-box nature. In contrast, non-LLM-based approaches, such as Brain (Yu et al., 2023) and ULP (Sedki et al., 2022), tend to be more interpretable, computationally efficient, and easier to deploy but lack the flexibility of LLMs in handling evolving log templates.
@@
LLM-based log parsing approaches typically require labeled datasets for fine-tuning or evaluation, though some methods leverage self-evolutionary approaches using previously parsed templates to substitute the need for a priori labels. However, with only $2$ and $4$ shots of labeled logs, LILAC (Jiang et al., 2024a) achieves superior performance on the corrected LogHub dataset (Khan et al., 2022) and Audit (Landauer et al., 2023b, 2021), compared to the conventional parsers and other LLM-based parsers, excluding LogBatcher (Xiao et al., 2024). Furthermore, LogBatcher demonstrates impressive performance on both runtime and effectiveness while being unsupervised and only requiring the log format as input parameter. Fine-tuning approaches also report outstanding performance, yet their stronger reliance on labeled logs and computational power and the robust and high performance of LogBatcher and LILAC demonstrate that ICL is a cheaper and performant alternative.
In comparison to the baseline parsers AEL (Jiang et al., 2008), SPELL (Du and Li, 2016), Drain (He et al., 2017), ULP (Sedki et al., 2022) and Brain (Yu et al., 2023), the LLM-based parsers of our evaluation selection require on average less configuration parameters. This finding is particularly noteworthy in light of the satisfactory performance achieved by LogBatcher (Xiao et al., 2024). This finding suggests that the deployment of LLMs can effectively substitute a substantial proportion of manual effort, thus enhance usability.
@@
It is evident that ICL and fine-tuning present powerful learning paradigms for LLM-based log parsing. They are not mutually exclusive to each other and can be employed simultaneously to improve parsing accuracy. ICL can be enhanced with RAG or with previously parsed or labeled logs as demonstrations. The ablative studies of reviewed papers (Jiang et al., 2024a; Xu et al., 2024; Xiao et al., 2024; Xu and Gau, 2024; Zhang et al., 2024a) unanimously conclude that dynamic demonstrations and sophisticated demonstration selection algorithms improve the accuracy scores. Smart caching solutions have been shown to lead to efficiency improvements, in terms of runtime and monetary expenses, since LLM calls are usually costly. Template revision methods that correct, delete or merge related templates retrospectively can further improve cached templates. This can also help adapt the parser to evolving logs due to updates or other behavioral changes of the computer systems without reconfiguration or retraining. Evaluations show that is also possible to improve the effectiveness of parsers with more powerful LLMs, yet the effective application of reasoning models like DeepSeek R1 should be investigated more closely. It has been shown that variable aware prompts improve the effectiveness in direct parsing with LLMs. Naturally, more explanations in the prompt mean more tokens per parsed log which negatively affects the inference time of LLMs. Short and concise instructions and explanations could thereby alleviate this issue. Furthermore, creating coarse-grained clusters to capture the commonalities of logs and fine-grained clusters to capture the variabilities of logs to support the template extraction process constitutes another promising technique, that can be applied by conventional and LLM-based parsers alike.
RQ4: What experiment design choices influence the comparability of the evaluation results between LLM-based log parsing approaches?
@@
Key experiment design choices include the selection of benchmark datasets, evaluation metrics, training methodologies, and hardware specifications. Consistency in log datasets and pre-processing methods is crucial for fair comparisons. For instance, not all parsers require the log format as an input parameter, such as LLM-TD (Vaarandi and Bahsi, 2024) and LUNAR (Huang et al., 2024). LUNAR explicitly includes log headers, such as IP addresses, timestamps, or levels, to grasp the full context of the log. For instance, partitioning logs into groups based on their timestamps is also possible with time intervals (Zhou et al., 2024). Previous research (He et al., 2016b) found that parsing only the extracted log message part, according to the log format, improves effectiveness, but especially LLMs with their generalization capabilities could utilize this extended context.
A set of metrics such as GA, PA, FTA, but also NED (or ED), should be standardized across studies to cover all relevant characteristics of templates, since each metrics maps to specific requirements of specific downstream tasks. The results of our evaluation show high variance between these metrics for different parsers, indicating that the selection covers a variety of aspects. The choice of model hyperparameters also significantly impacts comparability. For example, the researchers of Lemur (Zhang et al., 2025) report near-perfect scores of $0.999$ and $0.996$ for FGA and GA on LogHub (Zhu et al., 2023), yet they require log format, regular expressions, and four other (hyperparameter-tuned) parameters for each dataset. This demonstrates the potential performance levels that can be achieved. However, the number of parameters required is impractical for real-world applications and a comparison to the performance of other parsers is not feasible.
@@
The aforementioned experiment design choices concern all types of parsers. Specific to LLM-based parsers, the employed LLM is also an important factor to consider regarding comparability of evaluation results. As demonstrated, CodeLlama, GPT-3.5, and DeepSeek R1 perform significantly different on the effectiveness scores. Additionally, variations in API versions, underlying model updates, and differences in temperature settings further contribute to comparability issues in performance across studies. This highlights the need for standardized benchmarking practices to ensure fair and reliable comparisons between different LLM-based parsing approaches.
@@
The accessibility and reproducibility of the presented LLM-based parsing methods are significantly hindered by issues related to code availability, documentation, and execution reliability. Many approaches do not provide public code, and even when code is available, it often lacks essential components, such as necessary scripts or instructions, making replication difficult. Several repositories suffer from missing files, incorrect implementations, or incomplete functionality that does not align with the descriptions in their respective papers. Furthermore, discrepancies in the code, such as missing scripts for crucial steps or incorrect default settings, complicate reproducibility. Even when modifications are made to ensure compatibility — such as fixing typos, updating libraries, or implementing minor adjustments to improve consistency — these changes highlight the necessity of external intervention to make the methods functional. The need for such corrections aligns with broader findings from Trisovic et al. (Trisovic et al., 2022) and Olzewski et al. (Olszewski et al., 2023), who observed that a majority of replication code repositories contain errors that prevent immediate execution. Olzewski et al. further found that only a small proportion of the running codebases also produce the claimed results of the related papers (Olszewski et al., 2023). Additionally, some parsers require substantial configuration efforts, further limiting accessibility. Whilst minor improvements, such as adding caching mechanisms to DivLog (Xu et al., 2024) or modifying LogPrompt’s (Liu et al., 2024) prompt size, enhance usability, they also introduce deviations from the original implementations, raising concerns about reproducibility. The discrepancy between our own evaluation results and the reported results raises further concerns about the validity of the publications’ evaluations but naturally also about our own evaluation. Future benchmarks are expected to address these concerns. Finally, while log parsers based on LLM show great potential, their current state significantly hinders accessibility and reliable reproduction without substantial external effort.

### v2

@@ 6.3. Evaluation Metrics
Following previous work (Jiang et al., 2024b) we report Group Accuracy (GA), Parsing Accuracy (PA), and F1-score of Template Accuracy (FTA) which are described in Sec. 4.2. GA and PA are the most used ones in the literature (cf. Sec. 5.11) and FTA is recommended by Joang et al. (Jiang et al., 2024b). Section 4.2 also described the Edit Distance (ED), which is the recommended parsing metric of Petrescu et al. (Petrescu et al., 2023), from which we compute the Normalized Edit Distance (NED) (Marzal and Vidal, 1993). From our selection, NED is only used by Xiao et al. (LogBatcher) (Xiao et al., 2024). Nevertheless, we report NED because normalization makes it a more intuitive indicator of effectiveness than ED, as it fits into the $[0,1]$ range of the other conventional metrics.
@@ 6.4.1. Settings
The evaluation was performed using three LLMs of different sizes (small, medium, large):
@@
    GPT-3.5121212https://platform.openai.com/docs/models/gpt-3.5-turbo; accessed 7-March-2025 (gpt-3.5-turbo-0125) is a widely known closed-source model, reaching outstanding performance in the earlier stages of the LLM hype. It was chosen because it is the most used one from all reviewed papers and it is considered to be medium sized (the size is actually undisclosed but a paper by researchers of Microsoft (Singh et al., 2023) state it is 20 billion parameters). The model is proprietary and was accessed via the OpenAI API.
@@
    DeepSeek R1131313https://github.com/deepseek-ai/DeepSeek-R1; accessed 25-February-2025 utilizes a novel reinforcement learning approach for model training. It has recently received wide public attention due to its outstanding performance, while being relatively small (but still large) compared to similarly performing models like OpenAI’s o1 model, in combination with its open-source release. This model is used to assess the potential gains in accuracy that can be achieved by utilizing one of the most effective models (671 billion parameters). DeepSeek R1 was called via API from TogetherAI [^14] https://www.together.ai/; accessed 14-March-2025.
The runtime evaluation with the LogHub-2.0 dataset (Jiang et al., 2024b) was run on an Ubuntu 24.04 LTS server with a 32-core AMD EPYC-Milan CPU and 64 GB RAM. With the computation on this hardware and LLM calls to the GPT-3.5 API. In general, we use GPT-3.5 as the default LLM for our evaluation since it is the most used one by the selected approaches.
@@ 6.4.2. Configuration and Code Changes
As mentioned in Sec. 5.13.1, there are some issues with some of the parsers’ code which had to be resolved for the benchmark. To ensure the comparability of the output of the parsers, other changes were also necessary to the code of each parser. These changes are kept as small as possible to not disturb the original design or interfere with performance, or at least as little as possible. To ensure model or API compatibility, we updated deprecated versions of the OpenAI (python) package to the newer working ones and we added the functionality to call the OpenAI, Ollama, and TogetherAI API if not given. To have consistent output format, we ensure that the template placeholder symbols are always <*>. Since the parsers LLM-TD (Vaarandi and Bahsi, 2024) and LogPrompt (Liu et al., 2024) do not only parse the log content but the entire log, we modify the input to these parsers to only parse the content as well. As found by He et al. (He et al., 2016b), this typically improves the parsing accuracy.
@@
Specific changes were made to DivLog (Xu et al., 2024), which is one of the earliest works on LLM-based log parsing. It does not feature a caching mechanism and therefore, calls the LLM for each log line. Since this is a significant burden in terms of runtime (and cost) and therefore also unrealistic for real-life adaptation, we added a simple cache consisting only of the set of already parsed templates. The cache returns the corresponding template in the event of a match instead of calling the LLM. This slightly affects the accuracy metrics due to improved output consistency. Additionally, we implemented a minor post-processing step where multiple consecutive wildcards were replaced by a single wildcard <*> for efficient matching, to resolve an issue with the generation of infinite consecutive wildcards for some of the logs. This measure slightly increases DivLog’s accuracy.
@@
LogPrompt (Liu et al., 2024) accepts a maximum prompt size parameter which is by default $3000$, corresponding to roughly $25$ logs parsed at once (as they state). Early experiments revealed that this performs extremely poor, hence we set the maximum prompt size to $1000$. This corresponds to roughly $5$ to $10$ logs per prompt, which is in the recommended range of simultaneously parsed logs by Xiao et al. (Xiao et al., 2024). If a template cannot be matched with a log it is again sent to the LLM. We limit the maximum number of re-prompts to $3$ to avoid infinite LLM calls (which occurred in early experiments). LogPrompt also required some code corrections, other than typos. For some datasets, LogPrompt runs into an infinite loop due to a faulty template extraction process from the LLM responses. This was fixed by computing the enumeration of the logs instead of extracting it from the prompt again. From the paper it is not clear whether they used ICL, CoT or the self-prompt paradigm for their parsing performance evaluation, yet the parsing functionality is only given for the self-prompt case.
@@ 6.4.3. Sampling for Supervised Parsers
From the selected parsers LILAC (Jiang et al., 2024a) and DivLog (Xu et al., 2024) are supervised parsers and require labeled logs to achieve competitive accuracy. In this study, we emphasize minimal human effort and therefore keep the number of available templates $n$ small, namely $n=2$ and $n=4$. Note that the Apache dataset from LogHub (Zhu et al., 2023) contains the smallest number of templates, which is only $6$. For LILAC and DivLog it has been shown that the performance increases with an increasing number of labeled logs available (Jiang et al., 2024a; Xu et al., 2024). This is obvious as the similarity search retrieves the exact template the target log belongs to with a higher likeliness. Multiple works from our selection (Jiang et al., 2024a; Xu et al., 2024) as well as (Le and Zhang, 2023b) state that having a diverse candidate set is crucial to reduce the risk of overfitting to a specific log template, which is why they create specific sampling algorithms to sample from labeled logs. However, this demands having a large proportion of templates at hand and may require users to label logs manually, thus may hinder real-life application (Huang et al., 2024).
@@ 7.1.1. Baseline Performance
The performance of the baseline parsers is visualized in Fig. 3. We can see that neither of the baseline parsers achieves satisfactory performance on FTA and PA, meaning that most templates are not matched exactly nor correct when compared token-by-token.
@@ 7.1.2. Performance of LLM-based Parsers
The evaluation results of the parsers with the three LLMs on LogHub (Zhu et al., 2023) are given in Fig. 4(a), Fig. 4(b), and Fig. 4(c). Note that we excluded LogPrompt (Liu et al., 2024) from the evaluation with DeepSeek R1 because the LLM was barely able to extract any templates, resulting in scores of roughly 00 in experimental runs. The LLM did not provide the output in the structured format that was requested. Therefore, the template extraction process, which is based on markers positioned directly before the extracted template, was not able to properly extract the template. This resulted in long inner monologues of DeepSeek R1 and maximum reprompts due to log-template mismatches which would have resulted in an unreasonable financial expense given the cost of the LLM and the poor performance.
@@
In the boxplots of Fig. 4(a), Fig. 4(b), and Fig. 4(c) we can see that the best performance on each metric and for each LLM is achieved by LogBatcher (Xiao et al., 2024), followed by LILAC (Jiang et al., 2024a) for $2$ and $4$ shots. LogBatcher and LILAC clearly outperform the conventional parsers, while the conventional parsers outperform the rest of the LLM-based parsers. In general, the performance of each LLM parser is rather stable across the different models. CodeLlama visibly performs worst, but it is also by far the smallest of the models with 7 billion parameters. Given the size difference of GPT-3.5 and DeepSeek R1, it is surprising that DeepSeek R1 does not outperform GPT-3.5. Furthermore, we report the exact numbers for the evaluation with GPT-3.5 in Table 5 to show how the parsers performed on each individual dataset.
Table 5. Performance of the selected parsers with GPT-3.5 on the corrected LogHub datasets including Audit. The best scores for each metric on each dataset are marked bold.
Dataset
@@ 7.1.3. LLM General Performance
Figure 5. Averaged performance of LLM-based parsers for CodeLlama, GPT-3.5 and DeepSeek R1.
@@ 7.1.5. Performance on the Audit Dataset
An analysis of the performance of all parsers on the Audit dataset reveals that only the supervised parser, LILAC (Jiang et al., 2024a), and to some extent also DivLog (Xu et al., 2024), are able to parse the logs effectively. Despite the fact that the parameters of the baseline parsers were hyperparameter tuned, they did not achieve satisfactory results. It appears that the intended templates’ format of the logs is not straightforward for conventional parsers, but neither for GPT-3.5 without supervised demonstrations. This finding underscores the efficacy of supervised parsers, demonstrating that a mere two templates are sufficient to attain a reasonable performance level, potentially even a single one demonstrating the preferred template style.
Figure 7. Performance of the selected LLM-based parsers with GPT-3.5 and the baseline on the Audit data.
@@ 7.2. Efficiency
Figure 8 visualizes the computation time and LLM invocation time of the parsers, Fig. 9 shows the number of LLM calls made. The invocation time for conventional parsers is, of course, zero. Outstanding runtime efficiency is achieved by the conventional parser ULP (Sedki et al., 2022) on both datasets. While LLM-TD is the second fastest and also calls the LLM the least times it is rather the opposite on the BGL dataset. As mentioned in Sec. 5.13.1 and 6.4.2 there were some implementation problems with SelfLog (Pei et al., 2024). They provided a faulty script for the efficiency evaluation which is why we ran the evaluation with the script they designed for the evaluation of the $2000$ log lines version of LogHub (Zhu et al., 2023). However, for both HDFS and BGL the processes got killed and are therefore not featured in the plots.
@@ 8.1. Answers to Research Questions
LLM-based log parsing approaches offer key advantages such as adaptability to diverse log formats, the ability to generalize across different system logs, and robustness to unseen log templates as demonstrated by the evaluation on the Audit dataset (Landauer et al., 2023b, 2021) in Sec. 7.1.5 and other work (Fu et al., 2022). Approaches like the unsupervised parser LogBatcher (Xiao et al., 2024) and the supervised parser LILAC (Jiang et al., 2024a) can handle complex and unstructured logs more effectively than conventional methods, as demonstrated in Sec. 7.1.1 and Sec. 7.1.2 and their corresponding publications (Jiang et al., 2024a; Xiao et al., 2024). They utilize both syntactic and semantic information within log data, outperforming methods that focus solely on syntactic (e.g. Drain (He et al., 2017), SPELL (Du and Li, 2016)) or semantic information (e.g. DivLog (Xu et al., 2024)). However, they also come with drawbacks. The usage of LLMs comes with a significant computational burden, implying the requirement of hardware capable of running LLMs locally or using external API services. Methods like caching can significantly reduce the number of LLM calls and therefore also the latency, but especially systems with a high number of templates or frequently changing templates (e.g. due to updates) may still use too many resources for certain use cases. The use of LLMs also introduces randomness into the process due to the randomness within the LLM’s output and possibly overthinking for reasoning models (Cuadron et al., 2025). In contrast, approaches that do not integrate language models into their process, such as Brain (Yu et al., 2023) and ULP (Sedki et al., 2022), have deterministic output, are computationally efficient, and are often easier to deploy but lack the flexibility of LLMs in handling diverse log templates.
@@
LLM-based log parsing approaches typically require labeled datasets for fine-tuning or evaluation, though some methods leverage self-evolutionary approaches using previously parsed templates to substitute the need for a priori labels. However, with only $2$ and $4$ shots of labeled logs, LILAC (Jiang et al., 2024a) achieves superior performance on the corrected LogHub dataset (Khan et al., 2022) and Audit (Landauer et al., 2023b, 2021), compared to the conventional parsers and other LLM-based parsers, excluding LogBatcher (Xiao et al., 2024) (Sec. 7). Furthermore, LogBatcher demonstrates impressive performance on both runtime and effectiveness while being unsupervised and only requiring the log format as input parameter. Fine-tuning (Karanjai et al., 2024; Ma et al., 2024a; Mehrabi et al., 2024; Pang et al., 2024; Zhong et al., 2024) or even pretraining approaches (Ji et al., 2024) also report outstanding performance, yet their stronger reliance on labeled logs and computational power and the robust and high performance of LogBatcher and LILAC demonstrate that ICL is a cheaper but still performant alternative.
Compared to the baseline parsers AEL (Jiang et al., 2008), SPELL (Du and Li, 2016), Drain (He et al., 2017), ULP (Sedki et al., 2022) and Brain (Yu et al., 2023), the LLM-based parsers of our evaluation selection require on average less configuration parameters (Sec. 5.5). This finding is particularly noteworthy in light of the satisfactory performance achieved by LogBatcher (Xiao et al., 2024). This finding suggests that the deployment of LLMs can effectively substitute a substantial proportion of manual effort, and thus enhance the usability.
@@
It is evident that ICL and fine-tuning present powerful learning paradigms for LLM-based log parsing. They are not mutually exclusive to each other and can be employed simultaneously to improve parsing accuracy (Sec. 5.4). ICL can be enhanced with RAG or with previously parsed or labeled logs as demonstrations. The ablative studies of reviewed papers (Jiang et al., 2024a; Xu et al., 2024; Xiao et al., 2024; Xu and Gau, 2024; Zhang et al., 2024a) unanimously conclude that dynamic demonstrations and sophisticated demonstration selection algorithms improve accuracy scores (Sec. 5.6). Smart caching solutions show efficiency improvements in terms of runtime but also monetary expenses since LLM calls are usually costly (Sec. 5.7). Template revision methods that correct, delete, or merge related templates in postprocessing can further improve cached templates. This can also help adapt the parser to evolving logs due to updates or other behavioral changes of the computer systems without reconfiguration or retraining (Sec. 5.9). Evaluations show that it is also possible to improve the effectiveness of parsers with more powerful LLMs, yet the effective application of reasoning models like DeepSeek R1 should be investigated more closely (Sec. 7). It has been shown that variable-aware prompts improve the effectiveness in direct parsing with LLMs. Naturally, more explanations in the prompt mean more tokens per parsed log, which negatively affects the inference time of LLMs. Short and concise instructions and explanations could thus alleviate this issue. Furthermore, creating coarse-grained clusters to capture the commonalities of logs and fine-grained clusters to capture the variabilities of logs to support the template extraction process constitutes another promising technique, that can be applied by conventional and LLM-based parsers alike (Sec. 5.8).
RQ4: Which experiment design choices hinder the comparability, and hence which guidelines should be adopted in terms of configuration, datasets used, evaluation metrics, and reporting to make results comparable?
@@
Consistency in preprocessing methods is crucial for fair comparisons. For instance, not all parsers require the log format (see definition in Sec. 2) as an input parameter, such as LLM-TD (Vaarandi and Bahsi, 2024) and LUNAR (Huang et al., 2024). LUNAR explicitly includes log headers, such as IP addresses, timestamps, or levels, to grasp the full context of the log. For instance, partitioning logs into groups based on their timestamps is also possible with time intervals (Zhou et al., 2024). Previous research (He et al., 2016b) found that parsing only the extracted log message part, according to the log format, improves effectiveness, but especially LLMs with their generalization capabilities could utilize this extended context (Huang et al., 2024). In general, the choice of configuration parameters significantly impacts comparability. For example, the researchers of Lemur (Zhang et al., 2025) report near-perfect scores of $0.999$ and $0.996$ for FGA and GA on LogHub (Zhu et al., 2023), yet they require log format, regular expressions, and four other hyperparameter-tuned parameters for each dataset. This demonstrates the potential performance levels that can technically be achieved. However, the number of parameters required is impractical for real-world applications and complicates meaningful performance comparisons.
The datasets related to LogHub (Zhu et al., 2023; Jiang et al., 2008; Khan et al., 2022) are used in most publications, namely $25$ out of $29$. LogHub and its versions are widely used benchmarks, enabling direct comparison with a large body of existing research and providing a standardized way to evaluate parsing performance. We therefore recommend using at least one of the LogHub versions for the sake of comparability. However, relying solely on LogHub datasets may not cover all possible log structures and scenarios. Incorporating custom or other open-source datasets reveals how well parsers generalize to different types of logs, identifies their strengths and weaknesses, and avoids overfitting to a single dataset (or dataset collection) (Sec. 5.10).
We recommend a set of metrics, namely GA, PA, FTA, and NED, to be standardized across studies to cover the relevant characteristics of templates, since each metrics maps to specific requirements of specific downstream tasks (Sec. 5.11). This is important to prevent single-sided metric selection that focuses on single characteristics. Using FGA over GA should also be considered if class imbalance is a concern, but if not, there is little difference whether GA or FGA is used as long as the aspect of grouping performance is covered. The results of our evaluation in Sec. 7 show high variance between these metrics for different parsers, indicating that this selection covers a variety of aspects.
@@
The aforementioned experiment design choices concern all types of log parsers. Specific to LLM-based parsers, the LLM employed is also an important factor to consider regarding the comparability of evaluation results. As demonstrated, CodeLlama, GPT-3.5, and DeepSeek R1 perform significantly different on the effectiveness scores (Sec. 7). Additionally, variations in API versions, underlying model updates, and differences in temperature settings further contribute to comparability issues in performance across studies.
@@
The accessibility and reproducibility of the presented LLM-based parsing methods are significantly hindered by issues related to code availability, documentation, and execution reliability (Sec. 5.13.1). Many approaches do not provide public code, and even when code is available, it often lacks essential components, such as necessary scripts or instructions, making replication difficult. Several repositories suffer from missing files, incorrect implementations, or incomplete functionality that does not align with the descriptions in their respective papers. Furthermore, discrepancies in the code, such as missing scripts for crucial steps or incorrect default settings, complicate reproducibility. Even when modifications are made to ensure compatibility — such as fixing typos, updating libraries, or implementing minor adjustments to improve consistency (Sec. 6.4.2) — these changes highlight the necessity of external intervention to make the methods functional. The need for such corrections aligns with broader findings from Trisovic et al. (Trisovic et al., 2022) and Olzewski et al. (Olszewski et al., 2023), who observed that a majority of replication code repositories contain errors that prevent immediate execution. Olzewski et al. further found that only a small proportion of the running codebases also produce the claimed results of the related papers (Olszewski et al., 2023). Additionally, some parsers require substantial configuration efforts, further limiting accessibility (Sec. 5.5). Whilst minor improvements, such as adding caching mechanisms to DivLog (Xu et al., 2024) or modifying LogPrompt’s (Liu et al., 2024) prompt size, enhance usability, they also introduce deviations from the original implementations, raising concerns about reproducibility. The discrepancy between our own evaluation results (Sec. 7) and the reported results raises further concerns about the validity of the publications’ evaluations but naturally also about our own evaluation. Future benchmarks are expected to address these concerns. Finally, while log parsers based on LLM show great potential, their implementations’ current state significantly hinders accessibility and reliable reproduction without substantial external effort.